# Integrating Reinforcement Learning and Model Predictive Control with Applications to Microgrids

### v1

@@ 4.3 Differences between methods based on supervised and reinforcement learning
To shed light on the fundamental differences between methods based on supervised learning and our approach, we explain the main differences in the learning setups. From the perspective of supervised learning approaches, the problem becomes a classification problem, i.e. the network aims to predict the optimal sub-action for each time step of the prediction horizon. Instead of relying on decoupled Q-values to obtain the sub-actions, SL directly approximates the mapping from the augmented state to the optimal sub-actions. Evidently, this comes at the cost of repeatedly solving the MILP (5) to optimality to create the training dataset. Moreover, in SL, training does not aim at minimizing the long-term system’s operation cost as in RL, rather it targets at minimizing the classification error. The difference in the objective of RL and SL creates a notable difference in the behavior of the control policies, which are best shown in the case study of the next section. For a brief description of the training process in the context of supervised learning, the reader is referred to Appendix B.
@@ 5 Case study: microgrid control
In this section, we evaluate the performance of our approach for microgrid operation optimization. The goal is to minimize the long-term operation cost, typically defined over several hours, of the microgrid system described in Figure 4 and Appendix A. At each time step, the microgrid operator can determine a set of discrete and continuous decision variables, shown in Table 1. Herein we consider the receding horizon control strategy laid out in Section 3. In this case, the microgrid operator has to determine a sequence of discrete and continuous decision variables over a finite prediction horizon. Then, the underlying optimization problem is formulated as an MILP (15). Hence, efficiently solving such an MILP to minimize the required computation time is at the core of our problem. In this context, we compare our RL-based approach against standard MPC for MLD systems, where the optimization problem (2) is solved to optimality with branch-and-bound, and against the supervised learning approach described in the last section. To keep the comparisons fair, we use the general-purpose solver GUROBI [31] for the corresponding LPs and MILPs. Besides, as performance metrics, we analyze the total operation cost, feasibility rate, and computation time.
5.1 Setup
In the case study, we aim to solve the economic power dispatch problem for a microgrid, i.e. to fulfill the power demand of the microgrid loads while minimizing the operation cost. The knowledge of the market energy prices and the forecasts of the power demand and power generation from renewable energy sources – solar and wind energy are considered – allows the operator to plan and optimally schedule the operation of the microgrid over a time horizon. To accomplish this goal, the microgrid operator has a set of discrete and continuous actions. The discrete control actions are the following: buying or selling from grid $\delta_{\mathrm{grid}}(k)$, charging or discharging the energy storage system $\delta_{\mathrm{b}}(k)$ and turning on or turning off the generators $\{\delta_{i}^{\mathrm{dis}}(k)\}_{i=1}^{N_{\mathrm{gen}}}$ at time step $k$. The continuous actions are the power exchange with the main grid $P_{\mathrm{grid}}(k)$, power exchange with the energy storage system $P_{\mathrm{b}}(k)$, and the power provided by the dispatchable generators $P_{i}^{\mathrm{dis}}(k)$. For a summary of the decision variables, see Table 1. In this setting, the decision variables are represented by the vectors $\epsilon_{\mathrm{c}}(k)$ and $\epsilon_{\mathrm{d}}(k)$, which are formed by the concatenation of all discrete and continuous decision variables, respectively, over the prediction horizon. Furthermore, consider the exogenous signal $\gamma(k)$ contains the forecasts over the prediction horizon for the: purchasing, sale, and production prices; power generation from the renewable energy sources; and power demand from the loads.
@@ 5.1 Setup
The parameters of the microgrid considered in this case study are shown in Table 3. For the experiments, the number of dispatchable generators is set to $N_{\mathrm{gen}}=3$. Hence, for a prediction horizon $N_{\mathrm{p}}$, there are $5\cdot N_{\mathrm{p}}$ binary variables. As there is a binary variable for the connection with the main grid and another one for the connection with the storage unit, amounting to $5$ binary variables per time step. Hence, even for a small number of generators, the number of possible action sequences ($2^{5\cdot N_{\mathrm{p}}}$) is potentially huge depending on the chosen prediction horizon. With a sampling time of $T_{\mathrm{s}}=30$ minutes, all the methods are simulated with a prediction horizon of $N_{\mathrm{p}}\in\{4,12,24,48\}$, corresponding to look-ahead periods of 2, 6, 1226122,\ 6,\ 122 , 6 , 12, and $24$ hours, respectively. The objective of these experiments with a varying prediction horizon is to assess how the effectiveness of the methods scales with the increase of the number of binary variables.
@@
The renewable energy generation and load profiles were taken from the actual operation of the Dutch main grid, which is publicly available at [32]. Since the microgrid considered in this case study does not have the capacity to meet the demand of the entire grid, the real power profiles were linearly down-scaled to adjust the values to our case study. This is the only pre-processing operation applied to these signals, so the shapes of the renewable energy generation and load profiles remain unchanged. The RL-based and SL-based algorithms are trained on one year of data (2022) and their performances are evaluated in a different year (2021) to assess the capacity of the approaches to generalize and adapt to a new setting. The price profiles – $c_{\mathrm{prod}},\ c_{\mathrm{buy}},\ c_{\mathrm{sell}}$ – are not publicly available and, consequently, they were synthetically generated by three different normal distributions which obey a basic principle: on average producing energy is cheaper than purchasing energy from the main grid and higher than selling energy to the main grid. Due to the randomness of the prices, which mimic real market conditions, the price order can change and the microgrid operator can take advantage of some occasions. For instance, when the production price is smaller than the sell price, it is profitable to generate power that exceeds the load demand and sell the power surplus to the main grid.
Prediction horizon $N_{\mathrm{p}}=4$ $N_{\mathrm{p}}=12$ $N_{\mathrm{p}}=24$ $N_{\mathrm{p}}=48$
Learning Approach SL RL SL RL SL RL SL RL
Optimality gap 0.08% 0.79% 0.13% 1.73% 0.08% 2.01% 0.06% 4.36 %
Infeasibility rate (1/1000) 30 0 31.7 1.11 76 5.62 110.3 22.9
Reduction in computation time 57.06% 56.61% 69.82% 63.99% 71.14% 36.56% 78.05% 55.92 %
Table 2: Simulation results for the reinforcement learning (RL) and supervised learning (SL) approaches for solving the mixed-integer linear program (5). The optimality gap is the relative distance between the objectives of the linear program (6) using the learning approaches and the objective of the mixed-integer program (5) solved to optimality with branch-and-bound. The infeasibility rate is the frequency at which the learning approaches propose discrete actions that cause linear program (6) to be infeasible. The reduction in computation time is relative to the solution time of the MILP (5).
Figure 7: Reduction in computation time for different prediction horizons for the reinforcement learning (RL) and supervised learning (SL) approaches in the case study.
In our case study, the dispatchable units are assumed to be identical and, consequently, a single production price is used for all of them. As a result, for optimization purposes, several discrete actions are equivalent. For instance, there is no difference in having only the first or only the third generator turned on. This allows the action space to be reduced from $322^{5}=$ 32 actions to $16$ actions, making the action space easier to explore. For the RL-based approach, Boltzmann exploration is used during training so that exploration occurs according to the decoupled Q-values. In this type of exploration, the probability of a given sub-action being chosen is proportional to the softmax of its decoupled Q-value. The policy used for exploration is described below
@@
For the SL and RL approaches, the neural network architecture consists of a single layer of an LSTM network, whose outputs are connected to a fully connected layer. The approaches are then trained with several sizes of the hidden state of the LSTM network in the set $\{64,128,256,512\}$. Next, we report the best-performing trained neural networks.
@@ 5.2 Simulation results
The simulation results for both the methods and different prediction horizons are shown in Table 2 and Figures 5, 6, and 7. The methods are compared using three performance indices namely optimality gap, infeasibility rate, and computation time. The optimality gap is the relative distance to the optimal solution computed by branch-and-bound, i.e.
@@
$\mathrm{reduction\ in\ computation\ time}=\left(1-\frac{T_{\mathrm{learning}}}{T_{\mathrm{optimal}}}\right)\times 100,$
@@
where $T_{\mathrm{learning}}$ is the inference time of the LSTM added to the solution time of the linear program (6) and $T_{\mathrm{optimal}}$ is the solution time of the mixed-integer program (5). It is important to note that $T_{\mathrm{learning}}$ and $T_{\mathrm{optimal}}$ exclude the model-building time, which is required by the solver, as this would skew the comparison.
The simulation experiments shown in Table 2 and Figure 7 display the ability of the learning approaches to significantly reduce the online computation time of the MPC approach. The variability in the reduction in computation time stems from hyperparameter optimization (grid-search) with regard to the size of the LSTM hidden state. The computational cost for inference of the neural network decreases for smaller sizes of the hidden state, so the reported results reflect the most compact and best-performing policies. In the experiments, the SL approach performs better with a relatively small LSTM network (hidden state size is equal to $128$), so the reduction in computation is consistently lower than that of the RL approach.
@@
The reduction in computation time comes at the price of a small optimality gap and some instances of infeasibility. Regarding feasibility, the proposed approach consistently has significantly lower infeasibility rates than that of the supervised learning method across all the considered prediction horizons. However, the supervised learning method has a considerably lower optimality gap than that of the proposed approach. This reveals a trade-off between both learning approaches in our case study: supervised learning encourages optimality and reinforcement learning promotes feasibility. Hence, the results suggest that the selection among these learning approaches ultimately depends on the relative importance of optimality and feasibility for a given application.
Concerning the scalability for an increasing prediction horizon, as seen in Figure 5, the performance of the SL method scales well for the optimality gap. Alternatively, Figure 6 reveals that the RL approach scales much better than the SL method regarding the infeasibility rate. At last, as shown in Figure 7, the reduction in computation time slightly increases for the SL method and remains at around the same level for the RL approach.
It is important to stress that our approach entirely avoids the solution of MILPs as training is only based on the solution of LPs and the MPC role lies solely in the optimization of continuous variables. Despite this apparent benefit, the training time of the RL approach was considerably longer than that of the SL approach. The structure of the underlying MILP is a key factor in determining the training time of SL since this is the main bottleneck in the training of the SL approach. Hence, the comparison between the computation time of SL and RL approaches is problem-specific.
@@ 5.3 Discussion
The results in the previous section show high potential for employing the proposed framework to efficiently control MLD systems, where a good balance between computational time, optimality, and feasibility is achieved. What follows are several points of discussion highlighting important features of the framework:
  1. 1.
@@
    Even though discrete solutions $\epsilon_{\mathrm{d}}(k)$ that lead to the infeasibility of the LP (6) are discouraged, i.e. this event has a negative reward during training, the agent still may output an infeasible action. Moreover, feasibility can be restored either by an auxiliary control law, e.g. locally producing as much energy as possible and selling/buying the necessary power from the main grid, or by requesting a different action from the agent, e.g. selecting the second-highest Q-value instead of the highest. In the worst case, the mixed-integer linear program (5) can be solved to optimality if there is sufficient time.
  2. 2.
    In an MDP setting, some actions may not be allowed or may lead to outcomes that are not possible or feasible, depending on the state. In several implementations of RL algorithms, it is common to rule out these actions during training and inference, i.e. checking for feasibility and removing illegal actions from the action space. This typically speeds up training since the updates of the Q-function are focused only on the feasible actions. During inference, checking for feasibility can also have a positive effect in lowering the infeasibility rate. In the system considered herein, the feasibility of the action $\epsilon_{\mathrm{d}}(k)$ of the optimization problem (5) is only known after the action is sent to the environment and the LP (6) is solved. Since most of the constraints of the microgrid system depend on the continuous decision variables $\epsilon_{\mathrm{c}}(k)$, which are determined after the LP solution, it is difficult to eliminate infeasible actions before they are sent to the environment. However, even in this case, some actions can preemptively be ruled out. For instance, suppose that for a given load, the microgrid operator must turn on a minimum number of generators, and then some discrete actions that are evidently infeasible can be preemptively ruled out. Thus, implementing such a feasibility check may facilitate learning and improve performance.
  3. 3.
@@
    It is important to stress that our approach is agnostic to the reinforcement learning algorithm that is employed. As the RL algorithm, we chose Deep Q-Learning for its simplicity. Nonetheless, numerous algorithms, e.g. Rainbow DQN, PPO, and A3C, have shown better performance than Deep Q-Learning in several standard RL benchmarks. Thus, there is a possibility for improvements in optimality and feasibility rate if more sophisticated RL algorithms are implemented.
  4. 4.
    In the context of the proposed control framework, a fundamental difference between reinforcement learning and supervised learning lies in their goals. Reinforcement learning aims to maximize the long-term reward, which is related to minimizing both the operation cost and infeasibility rate. Supervised learning targets at solving a classification problem, i.e. at finding a mapping from states to optimal solutions. Intrinsically, reinforcement learning explores and learns about infeasible actions, while supervised learning narrowly focuses on predicting the optimal action. As a result, the output of reinforcement learning is richer in providing more information about the actions and their feasibilities. On the other hand, supervised learning is much more effective at its intended purpose of predicting the optimal solution at the cost of a higher infeasibility rate.
  5. 5.
    The microgrid model used in this paper has several simplifications when compared to a more realistic one as described in [1]. The implementation of the RL-based method proposed herein in more complex systems might reveal more about its advantages and drawbacks. Moreover, due to the similarity of the control problems, our approach has the potential to be applied to different energy systems and to other meaningful infrastructure systems such as transportation systems.
  6. 6.
    Although the proposed approach was used for the economic dispatch problem of microgrids, we highlight that it can be applied to any problem of the form (5), i.e. parametric mixed-integer quadratic (linear) problems that need to be solved repeatedly. Furthermore, our approach is easily extendable to other classes of problems, such as parametric mixed-integer nonlinear programs.
6 Conclusions and future work
This paper has presented a novel approach that combines reinforcement learning and model predictive control for mixed-logical dynamical systems. The main motivation for embedding learning into the optimization of the MPC controller is to reduce its online computational complexity. The core of our approach is to decouple the decisions on the discrete and continuous optimization variables, where the RL agent is responsible for computing the discrete variables and MPC optimizes over the continuous variables. Once the discrete variables are determined by RL, the MPC control problem is simplified from a mixed-integer quadratic (linear) program into a quadratic (linear) program, which significantly improves the required solution time. The approach has been implemented to solve the economic dispatch problem for a microgrid system and it has been compared against a branch-and-bound solver and a method that uses supervised learning to decouple the decision variables. Simulation results have shown that the RL agent, whose training is based on decoupled Q-functions, outputs policies that are close to optimality, showing that the decoupling does not hinder learning. Moreover, the computational time has been substantially decreased by the learning methods relative to the branch-and-bound solver. A comparison of the reinforcement learning and supervised learning techniques has revealed an interesting trade-off between the optimality gap and the infeasibility rate of the simplified linear program. While supervised learning favors performance, reinforcement learning outperforms in terms of feasibility.
The proposed approach is particularly targeted at reducing the computational burden relative to the growth of discrete decision variables caused by the increase of the prediction horizon. However, the complexity regarding the number of actions per time step is not addressed in our work. In this context, a topic of future work will be the investigation of methods to simplify the problem at each time step, to mitigate the overall combinatorial growth of the sequence of actions. Moreover, future work will also explore methods to further reduce the computation time, the optimality gap, and the infeasibility rate. Some possible directions to improve the performance have already been discussed in the previous section. The proposed framework can also be extended to control problems that require the solution of mixed-integer nonlinear programs, which paves the way for its application in even more complex systems.
CRediT authorship contribution statement
Caio Fabio Oliveira da Silva: Conceptualization, Methodology, Software, Validation, Investigation, Data Curation, Visualization, Writing – original draft, Writing – review & editing. Azita Dabiri: Conceptualization, Methodology, Supervision, Writing – review & editing. Bart De Schutter: Conceptualization, Methodology, Supervision, Writing – review & editing, Project administration, Funding acquisition.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v2

@@ 4.3 Differences between methods based on supervised and reinforcement learning
To shed light on the fundamental differences between methods based on supervised learning (SL), such as [12], and our approach, we explain the key differences in the learning setups. From the perspective of supervised learning (SL), the problem is formulated as a classification task, where the network predicts the optimal sub-action at each time step within the prediction horizon. Rather than relying on decoupled Q-values to derive sub-actions, SL directly approximates the mapping from the augmented state to the corresponding optimal sub-actions. This approach requires repeatedly solving the MILP (6) to optimality in order to construct the training dataset, which can be computationally expensive. Furthermore, unlike reinforcement learning, where the objective is to minimize the long-term operational cost of the system, SL focuses solely on minimizing the classification error during training. For a brief description of the training process in the context of supervised learning, the reader is referred to Appendix 8.
@@ 5 CASE STUDY: MICROGRID CONTROL
In the case study, we aim to solve the economic power dispatch problem for a microgrid system, i.e., to fulfill the power demand of the microgrid loads while minimizing the operation cost. The system is described in Appendix 7 and depicted in Fig. 4. The knowledge of the market energy prices and the forecasts of the power demand and power generation from renewable energy sources – solar and wind energy are considered – allows the operator to plan and optimally schedule the operation of the microgrid over a time horizon. To accomplish this goal, the microgrid operator has to determine a set of discrete and continuous actions over a finite prediction horizon. The discrete control actions are the following: buying or selling from grid $\delta_{\mathrm{grid}}(k)$, charging or discharging the energy storage system $\delta_{\mathrm{b}}(k)$ and turning on or turning off the generators $\{\delta_{i}^{\mathrm{dis}}(k)\}_{i=1}^{N_{\mathrm{gen}}}$ at time step $k$. The continuous actions are the power exchange with the main grid $P_{\mathrm{grid}}(k)$, power exchange with the energy storage system $P_{\mathrm{b}}(k)$, and the power provided by the dispatchable generators $P_{i}^{\mathrm{dis}}(k)$. For a summary of the decision variables, see Table 1. In this setting, the decision variables are represented by the vectors $\epsilon_{\mathrm{c}}(k)$ and $\epsilon_{\mathrm{d}}(k)$, which are formed by the concatenation of all discrete and continuous decision variables, respectively, over the prediction horizon. Furthermore, the exogenous signal $\gamma(k)$ represents the forecasts over the prediction horizon for the: purchasing, sale, and production prices; power generation from the renewable energy sources; and power demand from the loads.
@@
We consider the receding horizon control strategy laid out in Section 3. Consequently, the control problem can be formulated as the MILP of (6). Efficiently solving such an optimization problem is at the core of the problem. In this context, we compare our RL-based approach against standard MPC for MLD systems, where the optimization problem (2) is solved to optimality with branch-and-bound, and against the supervised learning (SL) proposed in [12] briefly described in the last section and in Appendix 8.
5.1 Setup
@@
The renewable energy generation and load profiles were taken from the actual operation of the Dutch main grid, which is publicly available at [37]. Since the microgrid considered in this case study does not have the capacity to meet the demand of the entire grid, the real power profiles were linearly downscaled to adjust the values to our case study. This is the only preprocessing operation applied to these signals, so the shapes of the renewable energy generation and load profiles remain unchanged. The RL-based and SL-based algorithms are trained on one year of data (2022) and their performances are evaluated in a different year (2021) to assess the capacity of the approaches to generalize and adapt to new data.
The price profiles – $c_{\mathrm{prod}},\ c_{\mathrm{buy}},\ c_{\mathrm{sell}}$ – are not publicly available, and, consequently, they were synthetically generated by three different normal distributions that obey a basic principle: on average, producing energy is cheaper than purchasing energy from the main grid and higher than selling energy to the main grid – examples of the cost profiles can be seen in Fig. 10 in Appendix 7. Due to the randomness of the prices, which mimic real market conditions, the price order can change, and the microgrid operator can take advantage of some occasions. For instance, when the production price is smaller than the sell price, it is profitable to generate power that exceeds the load demand and sell the power surplus to the main grid. The dispatchable units are assumed to be identical, and, consequently, the same production price is used for all of them.
Furthermore, the microgrid model, which is described in Appendix 7, is used for the training of the supervised learning and reinforcement learning agents, for the MPC controller, and for the system simulation. This ensures that performance differences arise solely from the algorithms. For the RL-based approach, Boltzmann exploration is used during training so that exploration occurs according to the decoupled Q-values. In this type of exploration, the probability of a given sub-action being chosen is proportional to the softmax of its decoupled Q-value. The policy used for exploration is described below
@@
For the SL and RL approaches, the neural network architecture consists of a single layer of an LSTM network, whose outputs are connected to a fully connected layer. The approaches are then trained with several sizes of the hidden state of the LSTM network in the set $\{64,128,256,512\}$. To simplify the presentation of the results, we only report the best-performing trained neural networks after hyperparameter tuning.
The methods are simulated with four prediction horizons $N_{\mathrm{p}}\in\{4,12,24,48\}$, corresponding to look-ahead periods of 2, 6, 1226122,\ 6,\ 122 , 6 , 12, and $24$ hours, respectively, given the sampling time of $T_{\mathrm{s}}=30$ minutes. The aim of these experiments is to assess how the effectiveness of the methods scales with the increase of the number of binary variables.
To ensure fair comparison, we use the general-purpose solver GUROBI [38] for the corresponding LPs and MILPs.
@@ 5.2 Simulation results
The simulation results for the methods and different prediction horizons are shown in Table 2 and Figs. 5, 6, and 7. As performance metrics, we analyze the optimality gap, infeasibility rate, and computation time. The optimality gap is the relative distance to the optimal solution computed by branch-and-bound, i.e.
@@
[…]
@@
where $T^{\mathrm{max}}_{\mathrm{i}}$ is the maximum of the inference time of the LSTM added to the solution time of the linear program (7) and $T^{\mathrm{max}}_{\mathrm{optimal}}$ is the maximum solution time of the mixed-integer program (6).
The simulation experiments, conducted over $3000$ randomly selected initial states and summarized in Table 2 demonstrate the effectiveness of learning-based approaches in significantly reducing the online computation time of MPC. In control applications, often the maximum computation is a more important criterion than the average computation time for hardware specification. As shown in in Fig. 7, which reports the reduction in the maximum computation time, the learning-based approaches are $8$ to $16$ faster than the branch-and-bound solver in the worst-case.
Optimality gap
Infeasibility rate
(1/1000)
Computation time
(mean/max/std)
$N_{\mathrm{p}}=4$ RL 0.20 1.6 1.2 / 2.1 / 0.1
SL 0.04 24.6 1.3 / 2.3 / 0.1
Optimal 0.00 0.0 3.4 / 19.7 / 0.7
$N_{\mathrm{p}}=12$ RL 0.36 2.3 2.0 / 3.6 / 0.2
SL 0.12 12.3 2.1 / 3.2 / 0.2
Optimal 0.00 0.0 8.0 / 46.5 / 2.3
$N_{\mathrm{p}}=24$ RL 0.59 1.0 4.8 / 9.4 / 0.4
SL 0.15 54.3 4.3 / 10.8 / 0.4
Optimal 0.00 0.0 23.6 / 84.8 / 3.3
$N_{\mathrm{p}}=48$ RL 0.76 11.0 10.9 / 14.5 / 0.5
SL 0.18 93.67 10.0 / 14.3 / 0.6
Optimal 0.00 0.0 116.2 / 231.3 / 8.5
Table 2: Simulation results for the reinforcement learning (RL), supervised learning (SL), and branch-and-bound approaches for solving the mixed-integer linear program (6). The optimality gap is the relative distance between the objectives of the linear program (7) using the learning approaches and the objective of the mixed-integer program (6) solved to optimality with branch-and-bound. The infeasibility rate is the frequency at which the learning approaches propose discrete actions that cause linear program (7) to be infeasible. The mean, maximum, and standard deviation of the computation time are reported.
[…]
44441212121224242424484848480050505050100100100100Prediction horizonInfeasibility rate (/1000$\cdot$/1000⋅ / 1000)RLSL
4444121212122424242448484848888810101010121212121414141416161616Prediction horizonReduction factorReduction in maximum computation timeRLSL
[…]
Figure 8: A box plot containing the computation time for the RL-based, SL-based, and MILP approaches. The median is highlighted in red. The blue box, only visible for the optimal (MILP) approach, extends from the first quartile to the fourth quartile, meaning that $50\%$ of the data lies in this region. The black points represent outliers. The black vertical lines represent the range of data, excluding outliers.
(a) Optimal solution (MILP).
(b) RL-based approach solution.
Figure 9: The figures above depict the solutions of the continuous variables for the optimal solution 9(a) and for the proposed approach 9(b) in a random time period of a $24$ hours selected from the testing dataset. The following continuous signals are shown: the power exchanged with the storage unit $P_{\mathrm{b}}$, the power exchanged between the microgrid and the main grid $P_{\mathrm{grid}}$, the total power generated by the microgrid $P_{\mathrm{dis}}$, the power generated by renewable energy sources $P_{\mathrm{res}}$, and the power demand of the microgrid $P_{\mathrm{load}}$.
@@
The reduction in computation time comes at the price of a small optimality gap – under $1\%$ – and some instances of infeasibility, represented in Fig. 6. The proposed approach consistently has significantly lower infeasibility rates than that of the supervised learning method across all the considered prediction horizons. The difference is more notable for $N_{\mathrm{p}}=48$, where the infeasibility rate for the proposed approach is $11/1000$ and that of the SL approach is $93.67/1000$. However, the supervised learning method has a slightly lower optimality gap than that of the proposed approach. This reveals a trade-off between both learning approaches in our case study: while supervised learning favors optimality, reinforcement learning outperforms in terms of feasibility. Hence, the results suggest that the selection among these learning approaches ultimately depends on the relative importance of optimality and feasibility for a given application.
A simulation of the power exchanges in the online operation of the microgrid is shown in Fig. 9, where both the behavior of the optimal controller and of the RL-based controller are given. Regarding performance, the optimality gap between the optimal solution (MILP) and the RL-based approach is $0.35\%$. The profiles of the continuous variables are almost identical for both the approaches, highlighting the RL-based method’s ability to closely replicate optimal operation. The associated costs for producing, selling, and buying energy are shown in Fig. 10 in Appendix 7. For brevity, discrete variables are omitted, as they can be directly inferred from the continuous signals, e.g., $\delta_{\mathrm{b}}(k)=1\iff P_{\mathrm{b}}(k)\geq 0$ and $\delta_{\mathrm{grid}}(k)=1\iff P_{\mathrm{grid}}(k)\geq 0$.
Concerning the scalability for an increasing prediction horizon, as seen in Fig. 5, the optimality of the proposed approach remains below $0.76\%$ and that of the SL-based approach remains under $0.18\%$, showing only small losses in optimality. The analysis of the infeasibility rate in Fig. 6 reveals that the RL approach scales considerably better than the SL method regarding the infeasibility rate. The contrast is especially evident at $N_{\mathrm{p}}=48$, where the infeasibility rate of the proposed approach is $8.5$ smaller. As shown in Fig. 7, the reduction in computation time is consistent across the prediction horizons, achieving its peak at $N_{\mathrm{p}}=48$ where the maximum computation time is reduced by a factor of $16$. Finally, a box plot for the computation time for all approaches is shown in Fig. 8.
@@ 5.3 Discussion
The experiments presented in the case study in the previous section demonstrate the potential of the proposed framework for efficiently controlling MLD systems, achieving a favorable balance between computational time, optimality, and feasibility. This section brings to discussion several important aspects of the proposed approach.
@@
Even though discrete actions $\epsilon_{\mathrm{d}}(k)$ that lead to the infeasibility of the LP (7) are discouraged, i.e., these actions have negative reward during training, the agent still may output an infeasible action. As feasibility is crucial in many applications, it can be restored either by an auxiliary control law, e.g., locally producing as much energy as possible and selling/buying the necessary power from the main grid, or by requesting a different action from the agent, e.g., selecting the second-highest Q-value instead of the highest. In the worst case, the mixed-integer linear program (6) can be solved to optimality if there is sufficient time. Therefore, it is straightforward to incorporate the aforementioned fallback mechanisms to ensure feasibility in online operation.
The sampling time in the case study is relatively large compared to the average solution time of the MILP. However, the proposed approach offers advantages beyond the average reduction in computation time. In control applications, it is crucial to ensure that the maximum solution time is predictable and remains within a reasonable bound, i.e., a solution must be computed before the next time step. For a standard mixed-integer linear program solver, the worst-case complexity can be large compared to the average solution time, since, in the worst-case, all the possible combinations of discrete variables must be enumerated and solved for. In contrast, the proposed approach only requires evaluating a recurrent neural network and solving a linear program, both of which exhibit predictable computational behavior. Hence, it is relatively easier to estimate a reasonable worst-case bound on the computation time needed to determine the control input for the proposed approach. Due to the relatively smaller worst-case bound, the proposed approach requires the allocation of a shorter time window for computing the control signal. As a consequence, the proposed approach can utilize more recent and accurate measurements of the system and exogenous signal, potentially improving the solution quality. Finally, for larger systems with more decision variables and possibly nonlinear objectives or constraints, the disparity in the worst-case computation times is expected to increase, as the complexity of the optimization problem grows at each node of the branch-and-bound tree.
@@
The proposed approach does not depend on a specific reinforcement learning (RL) algorithm. In this work, Deep Q-Learning was selected for its simplicity. Nonetheless, numerous algorithms, such as Rainbow DQN [39] and PPO [40], have shown better performance than that of Deep Q-Learning across several RL benchmarks. Therefore, it is likely that both optimality and feasibility rates could be further improved by employing more advanced RL algorithms.
A fundamental distinction between reinforcement learning and supervised learning lies in their objectives. Reinforcement learning seeks to maximize long-term rewards, which can be designed to minimize both operational costs and infeasibility rates. In contrast, supervised learning formulates a classification problem, aiming to establish a mapping from system states to optimal solutions. By its nature, reinforcement learning explores and evaluates infeasible actions, gaining insight into their consequences, whereas supervised learning focuses exclusively on predicting the optimal action. In the case study, the results suggest that reinforcement learning provides a richer output, offering more comprehensive information about actions and their feasibility. On the other hand, supervised learning is more accurate in predicting the optimal solution, but at the expense of a higher infeasibility rate. Further investigation is required to determine whether this observation extends to other problem domains.
In this work, modeling errors have not been considered. Nonetheless, the learning-based algorithms can be made more robust to model mismatch with techniques such as domain randomization [41] and dataset augmentation via noise injection in the system state [42].
The benefits of the learning-based methods are evident during online operation. However, training these methods requires accurate models and offline computational resources. For supervised learning, a dataset with the optimal solution of the system must first be obtained, after which the classifier weights are trained. In contrast, for reinforcement learning, the agent repeatedly interacts with a simulated system to obtain data to simultaneously adjust the decoupled Q-function approximator – possibly also a policy approximator – via temporal-difference learning. Fundamentally, with regard to training, the supervised learning approach requires the solution of mixed-integer programs, while its reinforcement learning counterpart requires the optimization of continuous decision variables. As a result, the computation cost in the reinforcement learning approach is smaller per iteration; however, it needs to perform more iterations due to exploration. Moreover, the determination of the best approach regarding offline computation load depends heavily on the type of available computational resources (e.g., CPU, GPU, and RAM) as well as the specific implementation of the data acquisition and training algorithms, particularly the extent to which the algorithm is designed to exploit existing hardware. Given these factors, the comparison of the computation load required by the different methods is complex and beyond the scope of this work.
6 CONCLUSIONS
This paper presented a novel approach that combines reinforcement learning and model predictive control for mixed-logical dynamical systems. The online computation time of the MPC controller can be significantly reduced by fixing the discrete decision variables with a policy trained by RL and then by optimizing over the continuous decision variables. This procedure effectively simplifies a mixed-integer linear program into a linear program.
For the training of the RL policy, we conceived the definition of the decoupled Q-function to decouple the discrete decision variables over the prediction horizon. Moreover, we showed how LSTM networks can be trained to approximate such a function.
Simulation experiments in a microgrid system revealed that our RL-based approach achieves a favorable trade-off between optimality, feasibility, and computation time. This suggests that the decoupled Q-function effectively captured the return of the sub-actions. Moreover, the computation time of the proposed approach was up to $16$ times faster than the standard MPC, whose solver was based on branch-and-bound.
Several potential research directions for enhancing the performance of the proposed approach were discussed in the previous section. Additionally, the framework can be readily extended to control problems involving mixed-integer nonlinear programs, enabling its application to even more complex systems.